OpenStudy (anonymous):

Would you make more interest by investing $3000 in a 6-month certificate of deposit (CD) at 3% interest or by investing $1000 in a 12 month CD at 3.5% interest? Show your work.

OpenStudy (phi):

Simple interest is I = P*r*t where P is principal, r is rate, and t is time The question does not say, so we must assume the interest is per year. the 6 month CD (1/2 of a year) gives I= 3000*0.03*0.5 the 12 month CD for 1 year gives I= 1000*0.035*1 (rate is 3.5%, principal is 1000, and time is 1 year)